If you invest $10,000 today and earn a 20% annual internal rate of return (IRR) over...

  1. If you invest $10,000 today and earn a 20% annual internal rate of return (IRR) over five years (with all of the proceeds received at the end of the fifth year), then the amount you will receive at the end of the fifth year is:

  1. How much would you pay today for an investment offering a lump sum of $100,000 in five years if you hoped to earn an annual rate of return of 25%?

  1. You invest $300,000 today and receive $800,000 cash proceeds at the end of five years. What is your internal rate of return on this investment? (to a tenth of a percent)
Answer #1

1). $10000 invested today at 20%

so after 5 year, amount received = PV*(1+r)^t = 10000*1.2^5 = $24883.20

2). amount received after 5 year = $100000 at rate = 25%

So its present value = FV/(1+r)^t = 100000/1.25^5 = $32768

3). invested today = $300000

amount received after 5 years = $800000

let interest rate be r

So, 800000 = 300000*(1+r)^5

=> 1+r = (8/3)^(1/5) = 1.2167

So, r = 0.2167 or 21.67%